E/D Night - UT 26-October-2020

1st half LBTO/Eng

2nd half LBTI Science

Night Center 00:004

-Mars observations will run 45 min over - Handover: 00:50

Night Plan

  • AO on-sky work
    • OSCO at twilight
    • Uranus Observatons (1hr)
      • /home/lbto/Service_Observing/Eng_D_Oct_25_31_2020/Mars_Uranus_Scripts
    • GRB 200826A - H-band Imaging with Bin 4 AO - (2hr)
      • /home/lbto/Service_Observing/Eng_D_Oct_25_31_2020/GRB_200826A
    • Mars Observations (1hr)
      • /home/lbto/Service_Observing/Eng_D_Oct_25_31_2020/Mars_Uranus_Scripts/
      • 2020-Oct-26 06:45 - 07:45 UT -local: 23:45-00:45 - first night is ok, although it spills over ~40minutes from night center
      • Plan for Observing:
        • Scripts manually edited to allow for H2+ND2
        • NSIGUI to Phobos (this is the AO Ref star)
        • Close the loop on Phobos
        • Use Telsvc for LUCI to offset to Mars ~15" away, absorb the offset
        • Continue script
    • ESM tests
      • Located in: /home/lbto/Service_Observing/Eng_D_Oct_25_31_2020/ESM (OBS-N)
      • M52 (sets 2nd half) AORef=10.7 and 15.2
      • M34 (rises above elevation 19:30hr) AORef = 12.4
  • Poor Weather Backup
    • LBC Pointing Model
      • /home/lbto/Service_Observing/PointingModel
    • Science Backup
      • /home/lbto/Service_Observing/Eng_D_Oct_25_31_2020/Poor_Weather
      • H-band Library (LUCI)
      • ICRF3 Followup (LUCI)

Calbration Files

  • For Mars observations: opticSieve mask and pupil images
    • For filenames: luci1.20201025.NNNN.fits

      0186 pupil image, aligned
      0188 dark

      I also took sieve mask images in K and 4 narrowband filters.

      0201 1x[1x60.0s] dark for NBs
      0202 1x[1x60.0s] PBeta
      0203 1x[1x60.0s] FeII
      0204 1x[1x60.0s] H2
      0205 1x[1x60.0s] BrG
      0206 1x[12x5.0s] K
      0207 1x[12x5.0s] dark for K
  • Night Log

    • Heavy cloud cover, particle count over limits, and winds 27-30m/s., We are unlikely to open to night.
    • Taking LUCI calibrations:
      • Darks 20x(1x2.5s) LIR NORM luci1.20201026.0001-20, luci2.20201026.0001-20
      • Darks 10x(1x20s) LIR NORM luci1.20201026.0021-30
      • Darks 10x(10x20s) LIR INT luci1.20201026.0031-40
      • Darks 10x(1x60s) MER INT luci1.20201026.0041-50
      • Darks 10x(6x10s) LIR INT luci1.20201026.0051-60
      • Darks 10x(26x2.51s) LIR INT luci1.20201026.0061-70, luci2.20201026.0021-30
      • Darks 10x(9x10s) LIR INT luci1.20201026.0071-80
      • Darks 10x(12x5s) LIR INT luci1.20201026.0081-90
      • Darks 10x(1x150s) MER NORM luci1.20201026.0091-100, luci2.20201026.0031-40
      • N30 Flats Pbeta luci1.20201026.0112-121
      • N30 Flats PGamma luci1.20201026.0122-131
      • Other N30 Flats were taken yesterday
    • Doug is taking over to run some NCPA measurements with ARGOS cal source
    • We verified remote desktop works on robs using solution in IT8196
    • UT5:00 Clouds have moved out, humidity is spiking, winds 30m/s, pariculates over limit
    • Light is only appearing in 2 out of 4 pupils and the PSF is very elongated. After much troubleshooting Doug has given up for today. Vibrations or something?
    • N30 Flats - lamp off/on
      • N30 Jlow luci1.20201026.0152-0161 5x(2x40) halo2
      • N30 Jhigh luci1.20201026.0162-0171 5x(2x37.5) halo2
      • N30 J luci1.20201026.0172-0181 5x(2x15.0) halo2
      • N30 OH1060 luci1.20201026.0182-0191 5x(3x21.0) halo1
      • N30 PGam luci1.20201026.0192-0201 5x(3x16.0) halo1
      • N30 PBeta luci1.20201026.0202-0211 5x(3x8.7) halo1
      • N30 OH1190 luci1.20201026.0212-0221 5x(3x13.0) halo1
      • N30 Y2 luci1.20201026.0222-0231 5x(3x3.0) halo1
      • N30 HeI luci1.20201026.0232-0241 5x(3x12.0) halo1
      • N30 FeII luci1.20201026.0242-0251 5x(3x4.5) halo1
      • N30 BrGam luci1.20201026.0252-0261 5x(2x5.5) halo1
      • N30 H luci1.20201026.0262-0271 5x(2x8.5) halo2
      • N30 H2 luci1.20201026.0272-0281 5x(2x4.5) halo1
      • N30 Ks luci1.20201026.0282-0291 5x(2x12.0) halo2
      • N30 K luci1.20201026.0292-0301 5x(2x8.5) halo2
    • Updated OT libraries with new times for N30 Luci1 and renamed to AO library rather than Diffraction Limited to reduce confusion. Uploaded.

-- BarryRothberg - 24 Oct 2020
Topic revision: r9 - 27 Oct 2020, JenniferPower
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback